Let f(x)=x^3+5x^2 and g(x)=2x^2-5.
f/g is undefined when x=
Note:Give your answer as a comma separated list.

Answer by stanbon(75887)

Let f(x)=x^3+5x^2 and g(x)=2x^2-5.
f/g is undefined when the denominator is zero.
Solve
2x^2-5 = 0
x^2 = 5/2
x = +sqrt(10)/2 , -sqrt(10)/2
--------------
Cheers,
Stan H.
--------------